OpenStudy (anonymous):

can someone check if my answer is correct, heres the problem, two identical positive charges has been separated by a 2.6 x 10^-2m, being released from rest. immediaately after the release, particle 1 has an accelelration a1 whose magnitude is 4.6 x 10^3m/s^2, while particle 2 has an acceleration a2 whose magintude is 8.5 x 10^3 m/s^2. find the unknown charge on each paticle and the mass of particle 2

OpenStudy (anonymous):

F = ma F = (6x10^-6)(4.6 x 10^3) F = 0.0276 N F = kq1q2/r^2 o.o276 = 9.0 x 10^9(x^2)/ (2.6x10^-2)^2 x = 5. 078 x 10^-8 mass1 m = f/a = 0.0276/8.5x10-3 m= 3.247x10^-6

OpenStudy (anonymous):

@KodyKthnx

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

yeah i have no idea.... xD

OpenStudy (anonymous):

@dan815 @alejandro1300

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@clynnew

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@Godlovesme

OpenStudy (godlovesme):

Sowwy no idea :/

OpenStudy (irishboy123):

post it on the physics forum centre of mass will be same throughout coulumbs law and newton laws of motion.